OpenStudy (anonymous):

Determine the rate law, including the values of the orders and rate law constant, for the following reaction using the experimental data provided. A + B products Trial [A] [B] Rate 1 0.15 M 0.12 M 1.3 × 10^-2 M/min 2 0.30 M 0.12 M 2.6 × 10^-2 M/min 3 0.15 M 0.24 M 5.2 × 10^-2 M/min

OpenStudy (anonymous):

here is what i have so far: rate= k[A]x[B]y trial 1/trial 2 : 1.3x10^-2/2.6x10^-2 = k[0.15]^x[0.12]^y/k[0.30]^x[0.12]^y 1.3x10^-2/2.6x10^-2 = [0.15]^x/[0.30]^x 1/2 = [1/2]^x x=1 trial 1/trial 3 : 1.3x10^-2/5.2x10^-2 = k[0.15]^x[0.12]^y/k[0.15]^x[0.24]^y 1.3x10^-2/5.2x10^-2 = [0.12]^y/[0.24]^y 1/4 = [1/2]^y y= 2 i just dont understand how to find the constant (k)